### v2

@@ Abstract
Deep neural networks often experience significant performance degradation when encountering domain shifts between training (source) and test (target) data. To mitigate this issue, Test Time Adaptation (TTA) methods have been introduced to adapt a pre-trained source model for handling out-of-distribution streaming target data. While these methods provide some improvement, they often lack a dependable mechanism for domain shift correction, which can be unpredictable in real-world applications. To address this challenge, we propose the Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that builds upon TTA by leveraging a few-shot support set. Following the principle of few inputs, big gains, FS-TTA effectively reduces the uncertainty of blind adaptation in previously unseen target domains. Furthermore, we introduce a two-stage framework to effectively tackle FS-TTA. The first stage involves fine-tuning the pre-trained source model using the few-shot support set while incorporating a feature diversity augmentation module to prevent overfitting. The second stage utilizes test time adaptation guided by a prototype memory bank, which facilitates the generation of high-quality pseudo-labels for model adaptation. Extensive experiments conducted on three cross-domain classification benchmarks validate the effectiveness, reliability, and superior performance of our proposed FS-TTA and framework.
keywords:
Test Time Adaptation, Few-Shot, Domain Shift, Transfer Learning.
Figure 1: Test Time Adaptation (TTA) vs. Few-Shot Test Time Adaptation (FS-TTA). FS-TTA incorporates a small number of labeled target samples, which can be easily collected offline before deployment with minimal annotation effort, in addition to the unlabeled target data used in TTA. The results for TTA are based on the performance of TENT (Wang et al, 2021) on the OfficeHome (Venkateswara et al, 2017).
1 Introduction
@@
In recent years, deep neural networks have exhibited remarkable capabilities in representation learning. However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al, 2015; Ganin and Lempitsky, 2015; Li et al, 2017). In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training. Besides, this distribution shift can significantly degrade the performance of the deployed source model.
@@
To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al, 2015; Tzeng et al, 2017; Zhang et al, 2019; Xiao and Zhang, 2021; Xin et al, 2023) and domain generalization (Volpi et al, 2018; Zhou et al, 2021; Kim et al, 2021; Sicilia et al, 2023). While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios. Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process. In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
@@
In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al, 2021) introduces fully test time adaptation (TTA). TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision. TTA is particularly focused on an online setting, where the model must adapt and make predictions immediately upon receiving each batch of potentially non-independent and identically distributed (non-i.i.d.) target samples. To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data. Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo, 2021; Wang et al, 2023), consistency regularization (Boudiaf et al, 2022), and anti-forgetting regularization (Niu et al, 2022). While these methods can perform model adaptation during the test time, they encounter three primary challenges:
@@
The fundamental reason for these challenges lies in the blind exploration of the target domain without any supervisory signal under domain shift. In practical scenarios, however, it is often feasible to obtain a small number of labeled target samples with minimal annotation cost. For instance, a few representative examples can be labeled offline before deployment by domain experts or system users. This leads us to ask: If given limited supervisory information from the target domain, could the adaptation performance be improved?
@@
Based on the aforementioned findings, we introduce the Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains. As illustrated in Figure 1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA. It is noted that such setting is particularly beneficial in scenarios where precision and reliability are paramount, such as medical image analysis and autonomous driving, even spending a few extra labeling costs.
To solve the FS-TTA, we develop an effective framework. For domain shift correction, we first fine-tune the pre-trained source model with the few-shot support set, fostering initial adaptation to the target. To prevent overfitting, we propose Feature Diversity Augmentation (FDA) to generate new features. During the test time, we employ a self-training strategy, which involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model online. Furthermore, in order to reduce the impact of noisy pseudo-labels on the model, we propose Entropy Filter and Consistency Filter. The former filters out high-entropy samples with low confidence, and the latter is achieved through dual-branch prediction consistency. The experimental results across various cross-domain image recognition datasets show that our FS-TTA method significantly surpasses the performance of state-of-the-art methods and other baselines. To sum up, our main contributions are as follows:
  • Emerging research direction: We highlight the setting of the Few-Shot Test Time Adaptation (FS-TTA), where a limited few-shot support set is available prior to test-time adaptation. Leveraging these few-shot samples enables more effective mitigation of domain shift.
  • Innovative framework: We propose a meticulously designed framework to address FS-TTA, incorporating fine-tuning of the pre-trained source model with a Feature Diversity Augmentation (FDA) module and performing test-time adaptation through high-quality pseudo-labeled samples in a self-training manner.
@@
Table 1: Comparison with various adaptation settings, where $s$ and $t$ denote source domain and target domain, respectively. $L^{d}$ and $U^{d}$ denote labeled datasets and unlabeled datasets from domain $d$. “Online” means that adaptation can predict a batch of incoming test samples immediately. “$k$” represents the number of samples per class. “C” indicates the number of classes for the target domain.
@@
Domain Generalization $L^{s_{1}},\dots,L^{s_{N}}$ 0
@@
[…]
2 Related Work
2.1 Domain Generalization
@@
Domain Generalization (DG) aims to train models on multiple related but distinct source domains to ensure effective performance on unseen target domains. To enhance robustness, DG techniques often employ strategies such as data augmentation (Huang et al, 2021; Volpi et al, 2018) and data generation (Zhou et al, 2021; Robey et al, 2021) to introduce greater diversity during training. Other prevalent approaches leverage representation learning to extract domain-invariant features. This includes kernel-based methods (Li et al, 2018b) that project data into a shared feature space, domain adversarial learning (Sicilia et al, 2023) that aligns distributions via adversarial objectives, and invariant risk minimization (Krueger et al, 2021) which encourages models to perform consistently across domains. In addition, self-supervised (Kim et al, 2021) and meta-learning-based techniques (Chen et al, 2022) have been explored to further improve generalization. However, without exposure to the target domain, generalization remains inherently limited.
2.2 Source-Free Domain Adaptation
@@
Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained source model to an unlabeled target domain while ensuring that no source data is accessed during the adaptation process. By eliminating the dependence on source data, SFDA effectively safeguards source data privacy, making it particularly suitable for scenarios where data sharing is restricted. SFDA techniques can be broadly categorized into two main approaches: pseudo-label strategies and generative methods. The former leverages target pseudo-labels to facilitate self-training, thereby enabling implicit adaptation without requiring explicit supervision (Tanwisuth et al, 2021; Ahmed et al, 2021; Liang et al, 2021; Xin et al, 2023). The latter employs generative models to synthesize target-style training data, allowing the model to bridge the domain gap through data augmentation and distribution alignment (Qiu et al, 2021; Liu et al, 2021). Similar to SFDA, our proposed Few-Shot Test Time Adaptation (FS-TTA) also maintains the source-free property, ensuring that adaptation is performed without relying on source data while leveraging a small support set to enhance adaptation efficiency.
2.3 Few-Shot Transfer Learning
@@
Test Time Adaptation (TTA) aims to adapt a pre-trained source model on-the-fly during inference to mitigate distribution shifts. Early TTA methods apply self-supervised learning objectives (Sun et al, 2020), but typically require access to training data or modification of the training process. TENT (Wang et al, 2021) addresses this by proposing fully test-time adaptation, relying solely on target samples and adapting batch normalization parameters via entropy minimization. Subsequent approaches such as (Schneider et al, 2020; Nado et al, 2020) update statistics on each incoming mini-batch, while methods like LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022) tackle catastrophic forgetting during continual adaptation. TSD (Wang et al, 2023) further integrates self-training to selectively update using confident predictions. As a result, adaptation often relies heavily on the quality of incoming test samples.
2.4 Test Time Adaptation
@@
Test Time Adaptation (TTA) aims to adapt a pre-trained source model during inference to mitigate distribution shifts between training and test domains. Early TTA methods address this challenge through self-supervised auxiliary tasks (Sun et al, 2020), which, while effective, often require access to training data or modifications to the training procedure. To overcome this limitation, TENT (Wang et al, 2021) proposes fully test-time adaptation by leveraging only target data, updating batch normalization parameters via entropy minimization. Building on this, subsequent works (Schneider et al, 2020; Nado et al, 2020) estimate batch normalization statistics dynamically from incoming test batches. Other approaches, such as LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022), focus on preventing catastrophic forgetting during continuous model updates. More recently, TSD (Wang et al, 2023) incorporates self-training by selecting confident test samples to guide adaptation. Despite their progress, these methods heavily rely on the quality and stability of online target data.
2.5 Comparisons with Other Settings
@@
We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Table 1.
@@
3 Preliminary
3.1 Instance Normalization
Instance Normalization (Ulyanov et al, 2016) is a normalization technique widely used in deep neural network architectures, especially in the context of style transfer and generative models. Let us consider a batch images with size $N\times C\times H\times W$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width of the images. For each sample $i$ and channel $c$, we compute the mean and standard deviation as follows:
@@ 3.2 Class Prototype
4 Method
4.1 Problem Setting
@@
4.2 Stage I: Model Adaptation via Fine-Tuning
To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set. Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process. To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics. Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model. This entire procedure is illustrated in Stage I of Figure 3.
Feature Diversity Augmentation (FDA). Prior research (Zhou et al, 2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision. To increase style diversity while preserving semantic consistency, we introduce Feature Diversity Augmentation (FDA), a feature-level data augmentation technique that simulates various image styles without altering the original class labels. This approach effectively enriches the support set and helps reduce the risk of overfitting during fine-tuning.
@@ 4.2 Stage I: Model Adaptation via Fine-Tuning
$\bm{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (5)
$\bm{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (6)
@@
$f_{i}^{\prime}=\bm{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\bm{\beta}_{\operatorname{mix}},$ (7)
@@
$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\bm{s}_{i}\right)\right),$ (8)
@@
4.3 Stage II: Test Time Adaptation
During this stage, a mini-batch of unlabeled samples, denoted as $x=\{x_{1},x_{2},..,x_{B}\}$, online arrives. The central concept of this stage is to employ a self-training strategy to update the fine-tuned source model online, enabling it to fully adapt to the target domain. This involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model. Thus, we first generate the pseudo-labels by $\hat{y_{i}}=\text{argmax}(p_{i})$ for $x_{i}$, where $p_{i}$ is the prediction logits. However, it is inevitable that there are always some noisy samples are misclassified, leading to wrong pseudo-labels. To address this issue, we propose two modules to produce high quality pseudo-labels. The first is entropy filter, which screens out unreliable samples using Shannon entropy (Shannon, 2001). Typically, samples with higher entropy are considered to have lower prediction confidence. The second module is a prototype memory bank classification, which works in tandem with the classifier. The prototype memory bank is used to generate pseudo-labels outside the classifier, according to the nearest class prototype in the feature space. After that, pseudo-labels with consistency prediction is preserved for model adaptation. The entire process is outlined in Stage II of Figure 3.
Entropy Filter. To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation. In this regard, we propose the Entropy Filter, which employs Shannon entropy (Shannon, 2001) to select confident samples in the mini-batch. For an sample $x_{i}$, its entropy can be computed as:

### v4

@@ Abstract
Deep neural networks often suffer severe performance degradation when deployed in target domains that differ from the training data. Test-Time Adaptation (TTA) seeks to alleviate this issue by updating a pre-trained model using only streaming unlabeled target data. However, the complete absence of supervision makes it difficult for TTA to correct misaligned decision boundaries, often resulting in unstable or even harmful updates under distribution shift. We propose Few-Shot Test-Time Adaptation (FS-TTA), a practical extension of TTA that leverages a small labeled support set from the target domain prior to deployment. Even minimal supervision substantially reduces the uncertainty of blind adaptation and provides the boundary information needed for more reliable model updates. Building on this insight, we introduce a unified two-stage adaptation framework. Stage I performs fine-tuning for boundary alignment using the labeled support samples, further enhanced by feature-level diversity to improve robustness in the low-shot regime. Stage II conducts test-time distribution refinement using the unlabeled target stream, where reliable pseudo-labels are produced via a prototype memory mechanism to ensure stable online adaptation. Extensive experiments on PACS, OfficeHome, and DomainNet demonstrate that the proposed FS-TTA setting and framework deliver consistently superior performance and significantly improve adaptation reliability over state-of-the-art TTA approaches.
I Introduction
@@
In recent years, deep neural networks have exhibited remarkable capabilities in representation learning. However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical [long2015learning, ganin2015unsupervised, li2017deeper]. In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training. Besides, this distribution shift can significantly degrade the performance of the deployed source model.
@@
To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation [long2015learning, tzeng2017adversarial, MDD, xiao2021dynamic, xin2023self] and domain generalization [volpi2018generalizing, zhou2021domain, kim2021selfreg, sicilia2023domain]. While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios. Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process. In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
Figure 1: Test Time Adaptation (TTA) vs. Few-Shot Test Time Adaptation (FS-TTA). FS-TTA incorporates a small number of labeled target samples, which can be easily collected offline before deployment with minimal annotation effort, in addition to the unlabeled target data used in TTA. The results for TTA are based on the performance of TENT [wang2020tent] on the OfficeHome [venkateswara2017deep].
@@
In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT [wang2020tent] introduces fully test time adaptation (TTA). TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision. TTA is particularly focused on an online setting, where the model must adapt and make predictions immediately upon receiving each batch of potentially non-independent and identically distributed (non-i.i.d.) target samples. To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data. Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling [IwasawaM21, wang2023feature], consistency regularization [boudiaf2022parameter], and anti-forgetting regularization [niu2022efficient]. While these methods can perform model adaptation during the test time, they encounter three primary challenges:
@@
The fundamental difficulty of fully unsupervised test-time adaptation is that the model must adjust to distribution shifts while blindly exploring the target domain without any reliable supervisory signal. This often results in unstable updates, making such methods difficult to deploy in real applications where robustness is essential. In many practical systems, it is both feasible and common to obtain a few labeled target samples before deployment, as domain experts or users can annotate representative examples during system setup. Even such minimal supervision can provide meaningful guidance during the initial stage of adaptation. This naturally leads to the question: If given limited supervisory information from the target domain, could the adaptation performance be improved?
@@
Building on this insight, we propose Few-Shot Test-Time Adaptation (FS-TTA), an adaptation framework that fully exploits the complementary roles of the labeled support set and the unlabeled test-time stream, encapsulating the principle of few inputs, big gains. The support set not only alleviates the initial domain shift, but also provides a reliable foundation for subsequent adaptation. In Stage I: Fine-tuning for Boundary Alignment, we use the labeled support samples to adjust the source model toward the target domain and reinforce this process with a Feature Diversity Augmentation (FDA) strategy, which regularizes the low-shot update and yields a more reliable initialization. On top of this initialization, Stage II: Test-Time Adaptation for Distribution Refinement, leverages the unlabeled test-time stream to further adapt the model. The broader distributional information contained in the stream complements the limited support set, while reliable pseudo-labels, selected according to confidence and consistency criteria, ensure that online updates remain stable and resistant to drift under domain shift.
  • Emerging research direction: We formalize Few-Shot Test-Time Adaptation (FS-TTA), a realistic setting where minimal target supervision is available prior to deployment. This setting bridges the gap between fully unsupervised TTA and traditional adaptation paradigms, highlighting the value of incorporating limited but reliable target-domain signals.
  • Innovative framework: We show that FS-TTA naturally decomposes into a unified framework consisting of boundary-level correction and distribution-level refinement, providing a coherent perspective on how labeled and unlabeled target-domain information should be jointly leveraged for stable and reliable adaptation.
@@
TABLE I: Comparison with various adaptation settings. where $s$ and $t$ denote source domain and target domain. $L^{d}$ and $U^{d}$ denote labeled datasets and unlabeled datasets from domain $d$. “Online” means that adaptation can predict a batch of incoming test samples immediately. “kk” represents the number of samples per class. “C” indicates the number of classes for the target domain.
@@
Domain Generalization $L^{s_{1}},\dots,L^{s_{N}}$ - 0
@@
Test Time Adaptation Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Mini-batch $U^{t}$ $|\text{mini-batch}|$, typically 128
Few-Shot Test Time Adaptation Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Few-shot support set $L^{spt}\subset L^{t}$ and mini-batch $U^{t}$ $k\times\text{C}$ and $|\text{mini-batch}|$
II Related Work
II-A Domain Generalization
@@
Domain Generalization (DG) aims to train models on multiple related but distinct source domains to ensure effective performance on unseen target domains. To enhance robustness, DG techniques often employ strategies such as data augmentation [huang2021fsdr, volpi2018generalizing] and data generation [zhou2021domain, robey2021model] to introduce greater diversity during training. Other prevalent approaches leverage representation learning to extract domain-invariant features. This includes kernel-based methods [li2018domain] that project data into a shared feature space, domain adversarial learning [sicilia2023domain] that aligns distributions via adversarial objectives, and invariant risk minimization [krueger2021out] which encourages models to perform consistently across domains. In addition, self-supervised [kim2021selfreg] and meta-learning-based techniques [chen2022discriminative] have been explored to further improve generalization. However, without exposure to the target domain, generalization remains inherently limited.
II-B Source-Free Domain Adaptation
@@
Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained source model to an unlabeled target domain while ensuring that no source data is accessed during the adaptation process. By eliminating the dependence on source data, SFDA effectively safeguards source data privacy, making it particularly suitable for scenarios where data sharing is restricted. SFDA techniques can be broadly categorized into two main approaches: pseudo-label strategies and generative methods. The former leverages target pseudo-labels to facilitate self-training, thereby enabling implicit adaptation without requiring explicit supervision [tanwisuth2021prototype, ahmed2021unsupervised, liang2021source, xin2023self]. The latter employs generative models to synthesize target-style training data, allowing the model to bridge the domain gap through data augmentation and distribution alignment [qiu2021source, liu2021source]. Similar to SFDA, our proposed Few-Shot Test Time Adaptation (FS-TTA) also maintains the source-free property, ensuring that adaptation is performed without relying on source data while leveraging a small support set to enhance adaptation efficiency.
II-C Few-Shot Transfer Learning
@@
Test Time Adaptation (TTA) aims to adapt a pre-trained source model on-the-fly during inference to mitigate distribution shifts. Early TTA methods apply self-supervised learning objectives [sun2020test], but typically require access to training data or modification of the training process. TENT [wang2020tent] addresses this by proposing fully test-time adaptation, relying solely on target samples and adapting batch normalization parameters via entropy minimization. Subsequent approaches such as [schneider2020improving, nado2020evaluating] update statistics on each incoming mini-batch, while methods like LAME [boudiaf2022parameter] and EATA [niu2022efficient] tackle catastrophic forgetting during continual adaptation. TSD [wang2023feature] further integrates self-training to selectively update using confident predictions. As a result, adaptation often relies heavily on the quality of incoming test samples.
II-D Test Time Adaptation
@@
Test Time Adaptation (TTA) aims to adapt a pre-trained source model during inference to mitigate distribution shifts between training and test domains. Early TTA methods address this challenge through self-supervised auxiliary tasks [sun2020test], which, while effective, often require access to training data or modifications to the training procedure. To overcome this limitation, TENT [wang2020tent] proposes fully test-time adaptation by leveraging only target data, updating batch normalization parameters via entropy minimization. Building on this, subsequent works [schneider2020improving, nado2020evaluating] estimate batch normalization statistics dynamically from incoming test batches. Other approaches, such as LAME [boudiaf2022parameter] and EATA [niu2022efficient], focus on preventing catastrophic forgetting during continuous model updates. More recently, TSD [wang2023feature] incorporates self-training by selecting confident test samples to guide adaptation. Despite their progress, these methods heavily rely on the quality and stability of online target data.
II-E Comparisons with Other Settings
@@
We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Table I.
@@
III Preliminary
III-A Instance Normalization
Instance Normalization [ulyanov2016instance] is a normalization technique widely used in deep neural network architectures, especially in the context of style transfer and generative models. Let us consider a batch images with size $N\times C\times H\times W$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width of the images. For each sample $i$ and channel $c$, we compute the mean and standard deviation as follows:
@@ III-B Class Prototype
IV Method
IV-A Problem Setting
@@
IV-B Stage I: Boundary Alignment via Few-Shot Fine-Tuning
A major challenge in test-time adaptation is that the source model’s decision boundary does not align with the target-domain distribution. The few-shot support set provides the only unbiased supervision capable of correcting this boundary misalignment.Therefore, Stage I aims to align the model boundary with the target domain by fine-tuning the pre-trained source model on these labeled samples. Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process. To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics. Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model. This entire procedure is illustrated in Stage I of Figure 3.
Feature Diversity Augmentation (FDA). Prior research [zhou2021domain] has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision. To increase style diversity while preserving semantic consistency, we introduce Feature Diversity Augmentation (FDA), a feature-level data augmentation technique that simulates various image styles without altering the original class labels. This approach effectively enriches the support set and helps reduce the risk of overfitting during fine-tuning.
@@ IV-B Stage I: Boundary Alignment via Few-Shot Fine-Tuning
$\boldsymbol{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (5)
$\boldsymbol{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (6)
@@
$f_{i}^{\prime}=\boldsymbol{\gamma}_{\operatorname{mix}}\odot\frac{f_{i}-\mu_{i}}{\sigma_{i}}+\boldsymbol{\beta}_{\operatorname{mix}},$ (7)
@@
$\mathcal{L}_{\text{cls}}=-\sum_{i=1}^{k*C}\mathcal{H}\left(y_{i},p\left(\hat{y_{i}}\mid\boldsymbol{s}_{i}\right)\right),$ (8)
@@
IV-C Stage II: Test Time Adaptation
After obtaining a target-aware initialization from Stage I, the model encounters a stream of unlabeled target samples $x=\{x_{1},x_{2},\ldots,x_{B}\}$ during deployment. These samples provide additional distributional insights beyond what the support set can offer. Stage II aims to refine the model online by leveraging this unlabeled stream while ensuring stable updates under distribution shift.
The central concept of this stage is to employ a self-training strategy to update the fine-tuned source model online, enabling it to fully adapt to the target domain. This involves assigning pseudo-labels to unlabeled online mini-batches and using these labels to further update the model. Thus, we first generate the pseudo-labels by $\hat{y_{i}}=\text{argmax}(p_{i})$ for $x_{i}$, where $p_{i}$ is the prediction logits. However, it is inevitable that there are always some noisy samples are misclassified, leading to wrong pseudo-labels. To address this issue, we propose two modules to produce high quality pseudo-labels. The first is entropy filter, which screens out unreliable samples using Shannon entropy [shannon2001mathematical]. Typically, samples with higher entropy are considered to have lower prediction confidence. The second module is a prototype memory bank classification, which works in tandem with the classifier. The prototype memory bank is used to generate pseudo-labels outside the classifier, according to the nearest class prototype in the feature space. After that, pseudo-labels with consistency prediction is preserved for model adaptation. The entire process is outlined in Stage II of Figure 3.
Entropy Filter. To dynamically update the model using online mini-batch target, it is crucial to filter out noisy samples, as they may be assigned to incorrect classes, resulting in inaccurate prototype computation. In this regard, we propose the Entropy Filter, which employs Shannon entropy [shannon2001mathematical] to select confident samples in the mini-batch. For an sample $x_{i}$, its entropy can be computed as: